OpenStudy (anonymous):

Larry can stuff 150 envelopes in one hour.Harold can stuff 225 envelopes in one hour.About how long will it take them to stuff 10,000 envelopes?

OpenStudy (anonymous):

150 + 225 = 375 envelopes in 1 hour. 10000/375 = 26.666666 hours (80/3 hours), which is also 1600 minutes.

OpenStudy (anonymous):

150 + 225 = 375 envelopes in 1 hour. 10000/375 = 26.666666 hours (80/3 hours), which is also 1600 minutes